a clinical trial tests a method designed to increase the probability of conceiving a girl. in the study 480 babies were born and 264 of them were girls. use the sample data to contruct a 99% confidence interval estimate of the percentage of girls born. based on the results, does the method appear to be effective?
Answer #1

p̂ = X / n = 264/480 = 0.55
p̂ ± Z(α/2) √( (p * q) / n)
0.55 ± Z(0.01/2) √( (0.55 * 0.45) / 480)
Z(α/2) = Z(0.01/2) = 2.576
Lower Limit = 0.55 - Z(0.01) √( (0.55 * 0.45) / 480) = 0.4915
upper Limit = 0.55 + Z(0.01) √( (0.55 * 0.45) / 480) = 0.6085
99% Confidence interval is ( 0.4915 , 0.6085 )
( 0.4915 < P < 0.6085 )

SInce the values in the confidence interval is positive, hence we can conclude that  method appear to be effective.